OpenStudy (anonymous):

2x - y = 1, if x = -1, then y = ? thank you<3

OpenStudy (usukidoll):

So your question is 2x - y = 1, if x = -1, then y = ? correct?

OpenStudy (usukidoll):

so we need to solve for y. Since we are given x = -1 you plug it into the equation and solve for y

OpenStudy (anonymous):

2x - (-1) = try this

OpenStudy (usukidoll):

so can you plug in x = -1 for this equation 2x - y = 1?

OpenStudy (anonymous):

the answer is -3 haha. thanks!

OpenStudy (usukidoll):

I'll check just to be sure 2(-1)-y =1 -2-y=1 +2 +2 -y=3 y= -3 yes you are correct :)

OpenStudy (usukidoll):

so x = -1 , y = -3 2(-1)-(-3) =1 -2+3=1 1=1 yup checks out :) good job!